OpenStudy (anonymous):

Please help me!! Will Give Medal!! Find f(5) for this sequence: f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

\[f(1)=2\\ f(2)=5\\ f(3)=f(1)+f(3)=2+5=7\] lather rinse repeat

OpenStudy (anonymous):

oops typo there, i meant \[f(3)=f(1)+f(2)=2+5=7\]

OpenStudy (anonymous):

Is that the answer or do I solve that? @satellite73

OpenStudy (anonymous):

Would the answer be 13 @satellite73

OpenStudy (anonymous):

@Tay111

OpenStudy (anonymous):

Unfortunately, i do not remember how to do this.

OpenStudy (anonymous):

Its Fine thanks Tho

OpenStudy (anonymous):

@amistre64 @Hero Can you help me please?

OpenStudy (amistre64):

you were given a rule to apply. what about the rule do you not understand? or is it the notation that is off putting?